According to the shop manual, when drilling a 11.8-mm-diameter hole in wood, plastic, or aluminum, a drill should have a speed of 1300 rev/min .

Part A

For a 11.8-mm-diameter drill bit turning at a constant 1300 rev/min , find the maximum linear speed of any part of the bit.

Part B

Find the maximum radial acceleration of any part of the bit.
Answer #1

diameter = 11.8 mm

r = d/2 = 0.0118 / 2 = 0.0059 m

maximum linear speed = r*w

                                = 0.0059*(2pi*1300/60)

                                = 0.8 m/s

B) maximum acceleration = v^2 / r

                      = 0.8^2 / 0.0059 = 1.1 m/s2

so maximum radial acceleration is 1.1 m/s2